# Neural Signals Generate Clinical Notes in the Wild

### 06 Mar 2026 v2

@@ Abstract
Generating clinical reports that summarize abnormal patterns, diagnostic findings, and clinical interpretations from long-term EEG recordings remains labor-intensive. We curate a large-scale clinical EEG dataset with 9,9229{,}922 reports paired with approximately 11,00011{,}000 hours of EEG recordings from 9,0489{,}048 patients. We therefore develop CELM, the first clinical EEG-to-Language foundation model capable of summarizing long-duration, variable-length EEG recordings and performing end-to-end clinical report generation at multiple scales, including recording description, background activity, epileptiform abnormalities, events/seizures, and impressions. Experimental results show that, with patient history supervision, our method achieves $70\%$$95\%$ average relative improvements in standard generation metrics (e.g., ROUGE-1 and METEOR) from 0.20.20.30.3 to 0.40.40.60.6. In the zero-shot setting without patient history, CELM attains generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. CELM integrates pretrained EEG foundation models with language models to enable scalable multimodal learning. We release our model and benchmark construction pipeline at https://github.com/Jathurshan0330/CELM.
Machine Learning, ICML
@@ 1 Introduction
Electroencephalography (EEG) records long-term, real-time neuronal activity with millisecond-level temporal resolution and is widely used in hospitals for neurological diagnosis, such as epilepsy (Kotoge et al., 2025) and sleep health (Yang et al., 2023). However, writing clinical reports to summarize diagnostic findings over long-term EEG recordings remains a labor-intensive process. Typically, neurologists visually inspect EEG signals to assess brain activity and then compose clinical reports that document abnormal patterns (i.e., EEG phenotypes) and their clinical interpretations (Tatum IV et al., 2016). Moreover, this task requires substantial domain knowledge and extensive clinical experience, which in turn incur long-term training costs and require continual updates as new EEG phenotypes or waveforms are identified (Biswal et al., 2019).
@@
To support this profound clinical practice, several automated clinical EEG report generation models based on advanced deep learning techniques have been proposed (Biswal et al., 2019; Lee et al., 2023; Yin and Shin, 2025). These methods aim to encode EEG recordings into latent representations that capture temporal structure and clinically relevant phenotypes, which are then decoded into clinical reports. However, existing methods remain limited in several aspects. ① Conceptually, some works formulate report generation as a phenotype classification or retrieval problem followed by text decoding. They do not model EEG-to-report generation as an end-to-end learning task, and the misalignment between classification and generation objectives does not guarantee optimal report generation. ② Methodologically, these methods operate on short EEG segments and fixed-context templates. They fail to model long-term temporal context and global diagnostic reasoning required for interpreting long-duration EEG data. Moreover, fixed-context templates prevent multi-granularity report generation beyond predefined categories. ③ Practically, these methods are developed as task-specific models, each tailored to a narrow reporting objective (e.g., impression-only reports) (Lee et al., 2023). In contrast, clinical practice requires multi-level reporting, where neurologists routinely construct overall EEG summaries and section-wise outputs, including impressions, interpretations, and event or seizure annotations.
Present work. We introduce CELM , which, to the best of our knowledge, is the first clinical EEG-to-language foundation model (ELM) capable of generating EEG reports at multiple scales. We train and evaluate CELM on 9,9229{,}922 clinical reports paired with approximately 11,00011{,}000 hours of EEG recordings from 9,0489{,}048 patients, spanning both clinical context–conditioned and zero-context settings, and covering key report sections including EEG description, background activity, epileptiform abnormalities, events/seizures, and impressions/interpretations.
Our motivation stems from recent advances in multimodal foundation models that enable text generation from non-text modalities (e.g., images and speech) (Bordes et al., 2024; Shen et al., 2025). Recent studies have begun to incorporate EEG signals into language-based modeling frameworks (Ndir et al., 2025; Gijsen and Ritter, 2025). However, they primarily leverage text supervision to enhance EEG classification and do not support EEG-to-text report generation.
To tackle aforementioned limitations, the contributions of this paper lie in: ❶ Conceptually, we formulate neurological clinical report generation as an EEG-to-Language foundation modeling problem. CELM enables end-to-end training and generation by directly translating raw EEG recordings into clinical text, without relying on intermediate phenotype classification or template-based pipelines. ❷ Methodologically, we introduce the first EEG-to-language multimodal framework, comprising three core components: (i) epoch-aggregated tokenization that produces compact representations from variable-length recordings, (ii) sequence-aware alignment that captures inter-epoch temporal dependencies, and (iii) prompt fusion for conditional report generation. Importantly, our model design harnesses the representation capabilities of pretrained EEG foundation models and LLMs while enabling flexible generation across multiple clinical settings. ❸ Practically, we introduce the first ELM benchmark construction pipeline based on the Harvard Electroencephalography Database (Sun et al., 2025; Zafar et al., 2025), which contains approximately $12\mathrm{k}$ clinical reports paired with about $100\mathrm{k}$ hours of EEG recordings from $10k$ patients. Rather than requiring access to the raw, unstructured 200200TB corpus and associated metadata, our pipeline enables efficient filtering of EEG recordings matched to reports. We release all source code to support further evaluation and development.
Experimental results show that CELM consistently outperforms all baselines across all settings. With patient history, it achieves $70\%$$95\%$ average relative improvements, with standard generation metrics (e.g., METEOR (Banerjee and Lavie, 2005) and ROUGE-1 (Lin, 2004)) from 0.20.20.30.3 to 0.40.40.60.6 across diverse cases. In the zero-context setting without patient history, CELM achieves generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. We provide multi-perspective case analyses to shed light on future ELM development.
@@ 2 Related Work
Clinical EEG Report Generation. An early attempt at EEG report generation, EEGtoText (Biswal et al., 2019), proposed a two-stage pipeline that first classifies EEG phenotypes, and then generates report text via a text decoder conditioned on the classified labels. However, this relies on intermediate phenotyping as a bottleneck, limiting the capabilities to capture nuanced clinical findings beyond predefined categories. While some attempts jointly learn EEG encoders and text decoders (Lee et al., 2023), they rely on fixed segmentation and template-based generation (Yin and Shin, 2025). As a result, these methods fall short of end-to-end clinical report generation from long-duration EEG recordings. We address the above challenges by introducing the first family of EEG–Language foundation models.
@@ 3 EEG-Report Benchmark Construction
Figure 1: Overview of our framework. (a) EEG–Report benchmark construction pipeline, including clinical report structuring (Section 3.1), matching reports to EEG sessions, and examples of standardized report sections. (b) The proposed Clinical EEG Language Model (CELM) comprises (b.1) Epoch-Aggregated Tokenization, (b.2) Sequence-Aware Alignment, and (b.3) Prompt Fusion and Generation.
@@
Currently, no publicly available benchmark provides structured EEG–report pairs for training and evaluating EEG-to-Language Models (ELMs). In this work, we leverage two resources from the Brain Data Science Platform: (1) the Harvard Electroencephalography Database v4.1 (Zafar et al., 2025; Sun et al., 2025), which contains approximately $10k$ clinical reports paired with about $100k$ hours of EEG recordings from 10,88610{,}886 patients across multiple hospital sites, and (2) the Electronic Health Records (EHR) Repository with corresponding unstructured neurology reports. The EHR repository links reports to one or more EEG sessions via temporal alignment between EDF start times and report timestamps. Therefore, we construct an EEG–report benchmark using a dedicated processing pipeline (Figure 1a). Specifically, Section 3.1 describes how neurology reports are extracted and organized based on whether their content is inferable from EEG data. Section 3.2 details the preprocessing and standardization of the matched EEG recordings.
3.1 Clinical Report Structuring
Neurology reports can be broadly decomposed into two types of sections: (1) Clinical Context, which contains information such as patient history and monitoring indications that cannot be inferred from EEG signals, and (2) EEG Findings, which describe EEG observations and clinical interpretations that are directly inferable from EEG recordings. This distinction is critical for evaluating clinical ELMs, as sections within EEG Findings provide the appropriate ground truth for benchmarking a model’s ability to generate clinical narratives from EEG signals.
@@ 3.1 Clinical Report Structuring
A key challenge lies in reliably extracting and structuring these sections without information loss. Substantial variability in reporting styles and formatting makes manual extraction infeasible at scale. We therefore propose a three-stage automated pipeline to extract and structure reports.
@@
Section Detection: first, we curate a comprehensive list of candidate section headers under each category and detect their presence using string matching.
@@
Section-wise Extraction: second, for each detected section, we employ a local LLM (Meta-Llama-3-8B-Instruct) to extract the corresponding text using simple copy-based prompts. This detect-then-segment strategy avoids hallucinations and cross-section contamination that arise when extracting multiple sections jointly. Local deployment is required, as data use agreements prohibit sending clinical reports to proprietary cloud-based LLMs.
@@
Canonical Normalization: finally, we standardize the extracted sections into canonical categories, including EEG description/details, impressions/interpretations, background activity, epileptiform abnormalities, and events/seizures.
3.2 EEG Preprocessing
@@
Matched EEG recordings are preprocessed using a standard pipeline. Signals are first band-pass filtered between 0.1–75 Hz and notch filtered at 60 Hz to remove power-line interference. The recordings are then resampled to 200 Hz to ensure compatibility with existing EEG-only foundation models. Each recordings are segmented into non-overlapping 10-second windows, which is the typical temporal context used by current EEG foundation architectures. Finally, EEG channels are standardized to a 22-channel montage following the International 10-20 system. While the dataset contains reports matched to multiple EEG sessions, we filter to samples with single-session matches for training, validation, and evaluation. We then perform patient-level splitting into 60/20/20 training, validation, and test sets to ensure no patient overlap across splits, thereby preventing data leakage and producing a structured EEG–report benchmark.
@@ 4 CELM: Clinical EEG Language Model
This section introduces our CELM , consisting of three key components: Epoch-Aggregated Tokenization, Sequence-Aware Alignment, and Prompt Fusion and Generation. Unlike classification-oriented EEG models that learn discriminative representations for categorical labels, clinical ELMs must identify and aggregate clinically relevant information from EEG recordings to support report generation at multiple levels. To this end, Section 4.1 discusses the unique challenges of clinical ELMs and provides an overview of the model architecture and forward process. Sections 4.2, 4.3, and 4.4 then present the details of the three components.
@@ 4.1 Architecture and Forward Process
Clinical EEG introduces three unique challenges that prevent the direct adoption of existing multimodal paradigms, such as dictionary-based alignment or contrastive learning with fixed-length epochs (Ndir et al., 2025). Here, we design CELM with three corresponding steps:
Epoch-Aggregated Tokenization addresses the extreme temporal scale of clinical EEG. EEG sessions span hours of high-frequency, multi-channel signals, resulting in at least $\sim 31.7\mathrm{M}+$ data points per recording, far exceeding the context limits of modern LLMs (Grattafiori et al., 2024; Team et al., 2025; Yang et al., 2025).
@@
Overall, given an input EEG recording, CELM processes the signal through above three stages. The recording is first transformed into compact epoch-level tokens via Epoch-Aggregated Tokenization, which reduces the extreme temporal scale. These tokens are then projected into the language model space through Sequence-Aware Alignment, enabling long-range temporal dependencies to be maintained across epochs. Finally, Prompt Fusion and Generation conditions the language model on the aligned EEG representations and optional clinical context to generate structured reports.
@@ 4.2 Epoch-Aggregated Tokenization
The fundamental bottleneck in developing ELMs is representing long EEG recordings within the LLM context constraints. A typical two-hour clinical EEG recording at $200$ Hz across $22$ channels yields $7,200\times 200\times 22\approx 31.7$M data points, far exceeding current LLM capacities, making raw sample-level tokenization infeasible. A common alternative is mini-window tokenization, leveraging pretrained EEG encoders that typically tokenize 1-second segments. However, naively applying this granularity to full recordings produces $\approx 158$K tokens per session, still well beyond the context limits of most local LLMs (Grattafiori et al., 2024; Team et al., 2025; Yang et al., 2025). To overcome this limitation while harnessing pretrained EEG representations, we propose Epoch-Aggregated Tokenization (see Figure 1b.1).
Given an EEG recording session $\mathbf{X}\in\mathbb{R}^{N\times C\times T}$, where $N,C$, and $T$ denote the number of epochs, channels, and time points per epoch (typically 10 seconds) respectively, we first tokenize mini $1$s windows within each epoch using a pretrained EEG encoder. The resulting representations are then aggregated across mini-windows and channels to produce a single token per epoch, yielding a compact sequence of epoch tokens $\mathbf{E}_{\text{eeg}}\in\mathbb{R}^{N\times D_{\text{eeg}}}$, where $D_{\text{eeg}}$ is the embedding dimension. The aggregation strategy is encoder-dependent. For example, CBraMod applies pooling (Wang et al., 2024b), whereas LaBraM (Jiang et al., 2024) and TFM-Tokenizer (Pradeepkumar et al., 2025) uses [CLS] pooling. This tokenization compresses the tokens by up to $C\times T$ ($\sim 220\times$ compared to mini-window tokenization), enabling EEG recordings to fit within LLM context limits.

### 13 May 2026 v3

@@ Abstract
Generating clinical reports that summarize abnormal patterns, diagnostic findings, and clinical interpretations from long-term EEG recordings remains labor-intensive. We present CELM, the first clinical EEG-to-Language foundation model capable of summarizing long-duration, variable-length EEG recordings and performing end-to-end clinical report generation at multiple scales. CELM integrates pretrained EEG foundation models with language models to enable scalable multimodal learning. We curate a large-scale clinical EEG dataset containing 9,922 reports paired with approximately 11,000 hours of EEG recordings from 9,048 patients to train CELM, and release the benchmark with an automated report-structuring pipeline to facilitate future research. Experimental results show that CELM consistently outperforms existing methods across all evaluation settings. Importantly, we further conduct human evaluation with clinical experts, demonstrating that CELM generates reports that are more clinically coherent, diagnostically reliable, and better aligned with expert interpretation. We release our model and benchmark construction pipeline at https://github.com/Jathurshan0330/CELM.
@@ 1 Introduction
Electroencephalography (EEG) records long-term, real-time neuronal activity with millisecond-level temporal resolution and is widely used for neurological diagnosis Yang et al. (2023); Kotoge et al. (2025). However, writing clinical reports to summarize diagnostic findings over EEG recordings remains a labor-intensive process. Typically, neurologists visually inspect EEG signals to assess brain activity and then compose clinical reports that document various phenotypes and their clinical interpretations (Tatum IV et al., 2016). This task requires substantial domain knowledge and extensive clinical experience, which incur long-term training costs and require continual updates as new EEG phenotypes or waveforms are identified Biswal et al. (2019).
@@
Limitations of prior works. Although several deep learning approaches have been proposed for EEG report generation Biswal et al. (2019); Lee et al. (2023); Yin and Shin (2025), several limitations remain. Conceptually, many existing methods formulate report generation as a phenotype classification task followed by text decoding. Such pipelines fail to model EEG-to-report generation in an end-to-end manner, leading to an objective misalignment between classification and generation. Methodologically, they rely on short EEG contexts (spanning only several to tens of seconds) and fixed-context templates that construct checklist-style mappings between EEG phenotypes and predefined clinical notes. They fail to model long-term reasoning required for interpreting long-duration EEG data, further preventing multi-granularity report generation. Practically, due to the limited resource, these methods are developed as task-specific models, each tailored to a narrow reporting objective Lee et al. (2023). However, clinical practice requires multi-level reporting, where neurologists routinely construct overall EEG summaries and section-wise outputs, including impressions, interpretations, and event or seizure annotations.
Figure 1: Overview of CELM, a clinical EEG-to-language foundation model (ELM) for translating long-duration EEG recordings into structured clinical reports. CELM addresses key challenges in EEG report generation, including limited EEG-report data, EEG-language alignment, and long-context EEG modeling. The framework constructs a large-scale EEG-report benchmark and enables direct generation of structured clinical reports from raw EEG recordings.
Present work. To fill the gap, we present CELM, to the best of our knowledge, the first Clinical EEG-to-language foundation model (ELM) that unlocks meaningful interpretation of raw clinical EEG signals and enable report generation at multiple scales. Our motivation stems from recent advances in multimodal foundation models that enable text generation from non-text modalities (e.g., images and speech) Bordes et al. (2024); Shen et al. (2025). CELM hence shifts the paradigm from conventional EEG classification toward language-centric EEG understanding and describing, with three key innovations: ❶ A report structuring pipeline that automatically decomposes clinical notes and generates hierarchical structured reports, spanning EEG description, background activity, epileptiform abnormalities, events/seizures, and impressions/interpretations. ❷ A novel pretrained foundation model that processes EEG recordings of up to three hours, learns multi-grained representations from local patterns to global temporal contexts, and translates representations into text for end-to-end report generation. This is achieved by three core components: (i) epoch-aggregated tokenization that produces compact representations from variable-length recordings, (ii) sequence-aware alignment that captures inter-epoch temporal dependencies, and (iii) prompt fusion for conditional report generation. Importantly, our model design harnesses the representation capabilities of pretrained EEG foundation models and LLMs while enabling flexible generation across multiple clinical settings. ❸ A large-scale EEG-report benchmark curated from the largest clinical EEG corpora to date (the Harvard Electroencephalography Database (Sun et al., 2025; Zafar et al., 2025)), comprising approximately 10,00010{,}000 clinical reports paired with over 11,00011{,}000 hours of EEG recordings from 10,00010{,}000 patients, enabling CELM training and facilitating future model development.
To rigorously evaluate CELM, we benchmark it against state-of-the-art (SOTA) methods across five real-world report drafting tasks. Moreover, we conduct human evaluation with six clinical experts to assess the quality and clinical reliability of the generated reports. Extensive experiments demonstrate that CELM not only achieves significant improvements across quantitative metrics, but also produces reports that are better aligned with human interpretation. Our contributions are as follows:
  • We introduce a novel EEG-to-text foundation model capable of learning representations from long-duration EEG recordings (up to three hours) and directly generating clinical reports.
  • We propose an automated report structuring pipeline for raw EEG-text clinical data, yielding 10,00010{,}000 clinical reports paired with over 11,00011{,}000 hours of EEG recordings from 10,00010{,}000 patients.
  • Beyond benchmarking CELM, we also conduct human evaluation to assess whether the generated reports are more diagnostically reliable. Results show that CELM is statistically significantly more clinically coherent than those produced by prior methods and general language models.
@@ 2 Related Work
Clinical EEG Report Generation. An early attempt at EEG report generation, EEGtoText (Biswal et al., 2019), proposed a two-stage pipeline that first classifies EEG phenotypes, and then generates report text via a text decoder conditioned on the classified labels. However, this relies on intermediate phenotyping as a bottleneck, limiting the capabilities to capture nuanced clinical findings beyond predefined categories. While some attempts jointly learn EEG encoders and text decoders Lee et al. (2023), they rely on fixed segmentation and template-based generation Yin and Shin (2025). As a result, these methods fall short of end-to-end clinical report generation from long-duration EEG recordings. We address the above challenges by introducing the first family of EEG–Language foundation models.
@@ 3 EEG-Report Benchmark Construction
Currently, no publicly available benchmark provides structured EEG–report pairs for training and evaluating EEG-to-text models. In this work, we leverage two resources from the Brain Data Science Platform: (1) the Harvard Electroencephalography Database v4.1 Zafar et al. (2025); Sun et al. (2025), which contains approximately $10k$ clinical reports paired with about $100k$ hours of EEG recordings from 10,88610{,}886 patients across multiple hospital sites, and (2) the Electronic Health Records (EHR) Repository with corresponding unstructured neurology reports. The EHR repository links reports to one or more EEG sessions via temporal alignment between EDF start times and report timestamps. Therefore, we construct an EEG–report benchmark using a dedicated processing pipeline (Figure 2).
Figure 2: EEG–Report benchmark construction pipeline, including clinical report structuring (Section 2), matching reports to EEG sessions, and examples of standardized report sections.
@@
Clinical Report Structuring. Neurology reports can be broadly decomposed into two types of sections: (1) Clinical Context, which contains information such as patient history and monitoring indications that cannot be inferred from EEG signals, and (2) EEG Findings, which describe EEG observations and clinical interpretations that are directly inferable from EEG recordings. A key challenge lies in reliably extracting and structuring these sections without information loss. Substantial variability in reporting styles and formatting makes manual extraction infeasible at scale. We therefore propose a three-stage automated pipeline to extract and structure reports.
@@
  • Section Detection: We curate a comprehensive list of candidate section headers under each category and detect their presence using string matching.
@@
  • Section-wise Extraction: For each detected section, we employ a local LLM (Meta-Llama-3-8B-Instruct) to extract the corresponding text using simple copy-based prompts. This detect-then-segment strategy avoids hallucinations and cross-section contamination that arise when extracting multiple sections jointly.
@@
  • Canonical Normalization: We standardize the extracted sections into canonical categories, including EEG description/details, impressions/interpretations, background activity, epileptiform abnormalities, and events/seizures.
@@
EEG Preprocessing. Matched EEG recordings are preprocessed using a standard pipeline. Signals are first band-pass filtered between 0.1–75 Hz and notch filtered at 60 Hz to remove power-line interference. The recordings are then resampled to 200 Hz to ensure compatibility with existing EEG-only foundation models. Each recordings are segmented into non-overlapping 10-second windows, which is the typical temporal context used by current EEG foundation architectures. Finally, EEG channels are standardized to a 22-channel montage following the International 10-20 system. While the dataset contains reports matched to multiple EEG sessions, we filter to samples with single-session matches for training, validation, and evaluation.
@@ 4 CELM: Clinical EEG Language Model
This section introduces our CELM, consisting of three key components: Epoch-Aggregated Tokenization, Sequence-Aware Alignment, and Prompt Fusion and Generation. An effective model must identify and aggregate clinically relevant information from long-term EEG recordings to support report generation at multiple levels. To this end, Section 4.1 discusses the unique challenges of clinical ELMs and provides an overview of the model architecture and forward process. Sections 4.2, 4.3, and 4.4 then present the details of the three components.
Figure 3: The proposed Clinical EEG Language Model (CELM) comprises (a) Epoch-Aggregated Tokenization, (b) Sequence-Aware Alignment, and (c) Prompt Fusion and Generation.
@@ 4.1 Architecture and Forward Process
Clinical EEG introduces three unique challenges that prevent the direct adoption of existing multimodal paradigms, such as dictionary-based alignment or contrastive learning with fixed-length epochs Ndir et al. (2025). Here, we design CELM with three corresponding steps:
Epoch-Aggregated Tokenization addresses the extreme temporal scale of clinical EEG. EEG sessions span hours of high-frequency, multi-channel signals, resulting in at least $\sim 31.7\mathrm{M}+$ data points per recording, far exceeding the context limits of modern LLMs Grattafiori et al. (2024); Team et al. (2025); Yang et al. (2025).
@@
Overall, given an input EEG recording, CELM processes the signal through above three stages. The recording is first transformed into compact epoch-level tokens via Epoch-Aggregated Tokenization, which reduces the temporal scale. These tokens are then projected into the language space through Sequence-Aware Alignment, enabling long-range temporal dependencies to be maintained across epochs. Finally, Prompt Fusion and Generation conditions the language model on the aligned EEG representations and optional clinical context to generate structured reports.
@@ 4.2 Epoch-Aggregated Tokenization
The fundamental bottleneck in developing ELMs is representing long EEG recordings within the LLM context constraints. A typical two-hour clinical EEG recording at $200$ Hz across $22$ channels yields $7,200\times 200\times 22\approx 31.7$M data points, far exceeding current LLM capacities, making raw sample-level tokenization infeasible. A common alternative is mini-window tokenization, leveraging pretrained EEG encoders that typically tokenize 1-second segments. However, naively applying this granularity to full recordings produces $\approx 158$K tokens per session, still well beyond the context limits of most local LLMs Grattafiori et al. (2024); Team et al. (2025); Yang et al. (2025). To overcome this limitation while harnessing pretrained EEG representations, we propose Epoch-Aggregated Tokenization (see Figure 3a).
Given an EEG recording session $\mathbf{X}\in\mathbb{R}^{N\times C\times T}$, where $N,C$, and $T$ denote the number of epochs, channels, and time points per epoch (typically 10 seconds) respectively, we first tokenize $1$s windows within each epoch using a pretrained EEG encoder. The resulting representations are then aggregated across windows and channels to produce a single token per epoch, yielding a compact sequence of epoch tokens $\mathbf{E}_{\text{eeg}}\in\mathbb{R}^{N\times D_{\text{eeg}}}$, where $D_{\text{eeg}}$ is the embedding dimension. The aggregation strategy is encoder-dependent. For example, CBraMod applies pooling Wang et al. (2024b), whereas TFM-Tokenizer Pradeepkumar et al. (2025) uses [CLS] pooling. This tokenization compresses the tokens by up to $C\times T$ ($\sim 220\times$ compared to window tokenization), enabling EEG recordings to fit within LLM context limits.